# ViSymRe: Vision Multimodal Symbolic Regression

### v3

@@ 6.10 Ablation analysis
We randomly select several equations from low-dimensional benchmarks and construct multi-scale test datasets by resampling coefficients and variable values ( see I for details). On these datasets, we evaluate the effectiveness of the proposed AMS algorithm in adapting ViSymRe to datasets with varying scales. We compare AMS with standard normalization techniques (Z-Score and Min-Max) and a baseline without scaling. As illustrated in Fig. 13, the AMS version demonstrates superior performance across all metrics. AMS is the only version capable of reliably recovering the ground-truth symbolic structure. In terms of Complexity, we can see that the AMS version generates equations with significantly lower Complexity compared to Z-Score and Min-Max versions, avoiding the overcomplexity problem caused by traditional scale scaling methods when recovering the symbolic structure of equations. Furthermore, the $R^{2}$ distribution and the Accuracy Solution Rate confirm that AMS version achieves robust numerical approximation results, achieving a 100% Accuracy Solution Rate.
@@
Fig. 14 illustrates the convergence results of the skeleton prediction loss ($\mathcal{L}_{vsp}$) for variants on the validation set. It can be observed that the Base model achieves the lowest loss. In contrast, the loss of the dataset-only variant is significantly higher than that of the others, and it stops converging as early as the 14th epoch. Furthermore, although the large Codebook exhibits a faster loss decrease in the early stages of training, its final convergence level is on par with the Base model, whereas the small Codebook, due to limited capacity, results in insufficient visual representation capability, leading to higher loss. However, overall, the Codebook size has a limited impact on the final convergence level of the model.
In addition, it is observed that, except for the dataset-only variant, the loss curves for all variants and the Base model exhibit a plateau lasting 3–4 epochs. For example, this phenomenon occurs during epochs 4–6 for the Base model and the variant with Codebook_size=256Codebook\_size=256, and during epochs 9–12 for the variant with Codebook_size=768Codebook\_size=768. This stagnation is attributable to the instability of the Codebook during the early stages of training.
Figure 14: Skeleton prediction loss of various ablation variant. The dataset-only variant exhibits suboptimal convergence and plateaus prematurely due to the absence of the visual modality. Furthermore, the impact of Codebook size on the final convergence level remains limited.
@@
Figure 15: Comparison of perplexity under different Codebook sizes Regardless of the initial Codebook size, the perplexity eventually converges to a similar level, demonstrating that the model is insensitive to the Codebook size.
@@
Table 7 presents the SR performance evaluation results of the Base model and its ablation variants on low-dimensional benchmarks. The results are the average of the median $R^{2}$ scores from 5 independent runs, along with their corresponding standard deviations. It can be observed that the performance ranking of the variants is generally consistent with the convergence levels reported in Fig. 14. The Base model achieves the highest $R^{2}$ score and exhibits the most stable performance. In contrast, the dataset-only variant exhibits the most pronounced degradation, with an average $R^{2}$ score dropping by more than 0.05 compared to the Base model, alongside increased fluctuations. Furthermore, the variant utilizing standard Cross-Attention yields a suboptimal solution, suggesting that the inadequate suppression of noise in the virtual vision, thus compromising the final SR performance.
@@ 7 Conclusions
Potential contributions to broader fields: In addition to the field of SR, the proposed architecture also has potential in a wider range of multimodal learning scenarios. Firstly, the Biased Cross-Attention mechanism offers a feasible solution for robust multimodal fusion in scenarios affected by noisy modalities (such as audio-visual recognition under noise). By dynamically suppressing task-irrelevant features via the bias term, it ensures stability even when one modality is unreliable. Secondly, the dual-visual pipeline architecture demonstrates how to effectively transform knowledge from data-rich training environments (such as multi-view sensors in robotics or multimodal medical records) into lightweight, resource-constrained inference environments. This module effectively bridges the gap in modality availability between training and deployment, providing a reference for efficient multimodal inference.
@@ Appendix E Supplementary analysis of AMS
In SR, normalization plays a different role compared to traditional machine learning tasks. While standard normalization methods (such as Z-score, Min-Max) aim to accelerate gradient descent, the primary objective in SR is to recover the ground-truth symbolic structure of equations. Furthermore, pre-trained SR models are typically trained on bounded data distributions. To enable these models to extrapolate to datasets with distributions of varying magnitudes, normalization must bridge the scale gap without corrupting the underlying symbolic form.To this end, we introduce AMS.
@@ Appendix F Supplementary results on Feynman benchmark
II.13.23 $\frac{x_{1}}{\sqrt{-\frac{x_{2}^{2}}{x_{3}^{2}}+1}}$ $x_{1}(0.237e^{\frac{2x_{2}^{2}}{x_{3}^{2}}}+0.764)$ $R^{2}\approx 0.9997$ 0
@@ Appendix G Supplementary results on ODE-Strogatz benchmark
d_bacres1 $\frac{x_{1}x_{2}}{0.5x_{1}^{2}+1}-x_{1}+20$ $\frac{1.195x_{1}^{1.077}e^{\frac{0.395\log{\left(x_{2}\right)}^{2.958}}{x_{1}^{1.455}}}}{\pi^{0.517}}+17.296$ 0.9999 0
@@
d_lv1 $x_{1}^{2}-2x_{1}x_{2}+3x_{1}$ $0.0596\cdot x_{1}\left(16.782\cdot x_{1}+\frac{33.564\cdot x_{2}}{(0.0493\cdot x_{1}+1)^{6.626\times 10^{-6}}}-50.345\right)$ 0.9999 0
d_lv2 $x_{1}x_{2}-x_{2}^{2}+2x_{2}$ $x_{2}\left(-1\cdot\frac{x_{1}}{(x_{1}+0.826548275822253)^{0}}-1\cdot x_{2}+2\right)$ 1.0000 1
@@ Appendix H Supplementary results on low-dimensional benchmarks
Appendix I Supplementary details on ablation analysis

### v4

@@ 6.10 Ablation analysis
To evaluate the effectiveness of the proposed AMS algorithm in adapting ViSymRe to datasets with varying scales, we randomly select several equations from low-dimensional benchmarks and construct multi-scale test datasets by resampling coefficients and variable values (see Table LABEL:tab:Supplementary_scale_ablation for details). We compare AMS with standard normalization techniques (Z-Score and Min-Max) and a baseline without scaling. As illustrated in Fig. 13, the AMS version demonstrates superior performance across all metrics. AMS is the only version capable of reliably recovering the ground-truth symbolic structure. In terms of Complexity, we can see that the AMS version generates equations with significantly lower Complexity compared to Z-Score and Min-Max versions, avoiding the overcomplexity problem caused by traditional scale scaling methods when recovering the symbolic structure of equations. Furthermore, the $R^{2}$ distribution and the Accuracy Solution Rate confirm that AMS version achieves robust numerical approximation results, achieving a 100% Accuracy Solution Rate.
@@
Fig. 14 illustrates the convergence results of the skeleton prediction loss ($\mathcal{L}_{vsp}$) for variants on the validation set. It can be observed that the loss of the dataset-only variant is significantly higher than that of the others, and it stops converging as early as the 14th epoch. Furthermore, although the large Codebook exhibits a faster loss decrease in the early stages of training, its final convergence level is on par with the Base model, whereas the small Codebook, due to limited capacity, results in insufficient visual representation capability, leading to higher loss. However, overall, the Codebook size has a limited impact on the final convergence level of the model.
In addition, it is observed that, except for the dataset-only variant, the loss curves for all variants and the Base model exhibit a plateau lasting 3–4 epochs. For example, this phenomenon occurs during epochs 4–6 for the Base model and the variant with Codebook_size=256Codebook\_size=256, and during epochs 9–12 for the variant with Codebook_size=768Codebook\_size=768. The possible reason for this stagnation is attributable to the instability of the Codebook during the early stages of training.
Figure 14: Skeleton prediction loss of various ablation variant. The dataset-only variant exhibits limited convergence and plateaus prematurely due to the absence of the visual modality. Furthermore, the codebook size exerts minimal influence on the ultimate convergence level.
@@
Figure 15: Comparison of perplexity under different Codebook sizes. Regardless of the initial Codebook size, the perplexity eventually converges to a similar level, demonstrating that the model is insensitive to the Codebook size.
@@
Table 6 presents the SR performance evaluation results of the Base model and main ablation variants on low-dimensional benchmarks. The results are the average of the median $R^{2}$ scores from 10 independent runs. It can be observed that the performance ranking of the variants is generally consistent with the convergence levels reported in Fig. 14. The Base model achieves the highest $R^{2}$ score and exhibits the most stable performance. In contrast, the dataset-only variant exhibits the most pronounced degradation, with an average $R^{2}$ score dropping by more than 0.05 compared to the Base model, alongside increased fluctuations. Furthermore, the variant utilizing standard Cross-Attention yields a suboptimal solution, suggesting that the inadequate suppression of noise in the virtual vision, thus compromising the final SR performance.
@@ 7 Conclusions
Potential contributions to broader fields: In addition to the field of SR, the proposed architecture also has potential in a wider range of multimodal learning scenarios. Firstly, the Biased Cross-Attention mechanism offers a feasible solution for robust multimodal fusion in scenarios affected by noisy modalities (such as audio-visual recognition under noise [71]). By dynamically suppressing task-irrelevant features via the bias term, it ensures stability even when one modality is unreliable. Secondly, the dual-visual pipeline architecture demonstrates how to effectively transform knowledge from data-rich training environments into lightweight, resource-constrained inference environments, such as multi-view sensors in robotics or multimodal medical records [69, 12, 33]. This module effectively bridges the gap in modality availability between training and deployment, providing a reference for efficient multimodal inference.
@@ Appendix E Supplementary analysis of AMS
In SR, normalization plays a different role compared to traditional machine learning tasks. While standard normalization methods (such as Z-score, Min-Max) aim to accelerate gradient descent, the primary objective in SR is to recover the ground-truth symbolic structure of equations. Furthermore, pre-trained SR models are typically trained on bounded data distributions. To enable these models to extrapolate to datasets with distributions of varying magnitudes, normalization must bridge the scale gap without corrupting the underlying symbolic form.
@@ Appendix F Supplementary results on Feynman benchmark
II.13.23 $\frac{x_{1}}{\sqrt{-\frac{x_{2}^{2}}{x_{3}^{2}}+1}}$ $x_{1}(0.237e^{\frac{2x_{2}^{2}}{x_{3}^{2}}}+0.764)$ $0.9997$ 0
@@ Appendix G Supplementary results on ODE-Strogatz benchmark
d_bacres1 $\frac{x_{1}x_{2}}{0.5x_{1}^{2}+1}-x_{1}+20$ $\frac{1.195x_{1}^{1.077}e^{\frac{0.395\log{\left(x_{2}\right)}^{2.958}}{x_{1}^{1.455}}}}{\pi^{0.517}}+17.296$ 0.9977 0
@@
d_lv1 $x_{1}^{2}-2x_{1}x_{2}+3x_{1}$ $0.06\cdot x_{1}\left(16.782\cdot x_{1}+\frac{33.564\cdot x_{2}}{(0.05\cdot x_{1}+1)^{6.626\times 10^{-6}}}-50.345\right)$ 0.9999 0
d_lv2 $x_{1}x_{2}-x_{2}^{2}+2x_{2}$ $x_{2}\left(-1\cdot\frac{x_{1}}{(x_{1}+0.827)^{0}}-1\cdot x_{2}+2\right)$ 1.0000 1
@@ Appendix H Supplementary results on low-dimensional benchmarks
Appendix I Supplementary results on Phenomenological & First-principles benchmark
Dataset equation $R^{2}$
bode $0.15e^{0.68x_{1}}+0.39$ 0.997
hubble $480x_{1}+169\sin{\left(10x_{1}+9.11\right)}+25$ 0.925
kepler $362x_{1}^{1.5}$ 0.999
ideal_gas $\log{\left(\frac{10.5x_{1}x_{2}}{x_{3}}-6.93\right)}$ 0.998
leavitt $0.43x_{1}+11.8+3.14e^{-0.32x_{1}^{3}}$ 0.980
rydberg $\frac{3.55x_{1}-3.73}{x_{2}^{0.69}}-15.9$ 0.993
absorption $0.0045x_{1}+3.57-3.43e^{-0.032x_{1}^{1.41}}$ 0.995
supernovae_zg $0.94-0.93e^{-\frac{89.4}{x_{1}^{2}}}$ 0.981
Appendix J Supplementary details on ablation analysis